OpenStudy (anonymous):

f(x)=(x+1)(x-7) transform it to vertex form and find its h and k

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) first factor it out what would you get?

OpenStudy (sleepyhead314):

then vertex form is y = a(x-h)^2 + k you can get this from the y = ax^2 + bx + c by taking the b value, dividing it by 2,2 then squaring it then, keeping the original b value, adding a number to the c to get it equal to the squared half of b, but when adding a number remember to subtract the same amount along with it like if you are adding a 18 then you should also subtract a 18 on the same side do you follow so far? or would you like an example? :)

OpenStudy (anonymous):

Well I'm already done actually with that one but thank you. You're better than my teacher. Her teaching is different but thank you though O:)

OpenStudy (sleepyhead314):

I'm glad I could help ^_^